OpenStudy (bbb911):

This is Geometry: The measure of the angles of a triangle are x + 2, x +4, and 2x + 10. Find the measure of the smallest angle.

OpenStudy (anonymous):

(x+2)+(x+4)+(2x+10)=180 4x+16=180 4x=180-16 4x=164 x=41 smallest angle=x+2=41+2 =43

OpenStudy (bbb911):

thank you oh so much